Give an indirect proof of the theorem; “If 𝑛 is an integer and 

           n3 +13is odd, then n is even.”?  

1
Expert's answer
2021-05-07T10:30:04-0400

Solution:

Indirect proof can be done as ‘proof by contraposition’.

Proof by contraposition:

The contraposition of the statement is "If n is odd then "n^{3}+13" is even,".

Hence, to proof the contraposition, we need to assume that n is odd.

By the definition of odd numbers, there is an integer k such that

"n=2 k+1"

On substituting "n=2 k+1" into "n^{3}+13" , we get

"n^{3}+13=(2 k+1)^{3}+13=\left(8 k^{3}+12 k^{2}+6 k+1\right)+13=8 k^{3}+12 k^{2}+6 k+14=2\left(4 k^{3}+6 k^{2}+3 k+7\right)"

Thus, we can find an integer "m=4 k^{3}+6 k^{2}+3 k+7" such that

"n^{3}+13=2 m"

It shows, "n^{3}+13" is even.

Since the contraposition is true then the original statement is also true.